can you show me how to divde 4 7/8 by 1/2?
Answer by sophxmai(62)

You want to divide 4 7/8 by 1/2

First make the 4 7/8 into an improper fraction
(4*8+7)/8=39/8

Now you have (39/8) divide by (1/2)
Flip the fraction after the divide sign and change the sign into a multiply sign

(39/8) x (2/1)

Now multiply them together
(39/8) x (2/1)
=(39/4)

If it says to make it into a proper fraction again, do so
39/4= 9 3/4